The volatile liquid ethyl mercaptan, C2H6S, is one of the most odoriferous substances known. It is sometimes added to natural gas to make gas leaks detectable. How many C2H6S molecules are contained in a 1.0 mL sample? The density of liquid ethyl mercaptan is 0.84 g/mL.


1
Expert's answer
2021-11-11T14:15:01-0500

Mass in 1ml= "density \times volume=0.84\times 1 = 0.84g"

Molar mass of "C_2H_6S=(12.011\times 2+1.008\times 6+32.06\times 1)g/mol"

Moles of "C_2H_6S=mass \/molar mass=0.84\/62.134=0.0135moles"

Since, 1 mole molecules contain "6.022\times 10^{23}molecules"

number of molecules present in 0.0135 moles "=0.0135\times 6.022\times 10^{23}=8.13\times 10^{21}molecules"